OpenStudy (anonymous):

4x^2-32x+48 _______________ 3x^2-17x-6

OpenStudy (anonymous):

(4 (x-2))/(3 x+1)

jhonyy9 (jhonyy9):

4(x2 -8x +12)=4(x-6)(x-2)

jhonyy9 (jhonyy9):

3x2 -17x -6= (3x+1)(x-6)

jhonyy9 (jhonyy9):

simplify by (x-6) and check what will remain 4(x-2)/(3x+1)